# Imaginaries, products and the adele ring

### v2

@@ 3.7. Definable groups and automorphism groups
To bridge the gap between Borel and closed, we will use a result of Becker and Kechris. Taking into account the Harrington-Kechris-Louveau theorem [10] or [1] 3.4.3, Theorem 3.4.5 of [1] asserts: Let $\mathbf{X}$} be a Polish space, $G$ a group acting on $\mathbf{X}$} by homeomorphisms, and let $R$ be the equivalence relation on $\mathbf{X}$} whose classes are the $\Gamma$-orbits. If there is a dense orbit and $R\subset{\mathbf{X}}^{2}$ is meager, then $R$ is not smooth.
@@
If $H$ is meager in $\mathbf{H}$}, it is easy to see that $R$ is meager in ${\mathbf{H}}^{2}$; but this contradicts [1] 3.4.5. Hence $H$ is not meager. By [18] Theorem 9.9, $H=HH^{-1}$ contains an open neighborhood $U$ of the identity of $\mathbf{H}$} ; hence $H=HU=\bigcup_{h\in H}hU$ is open in $\mathbf{H}$}, and thus also closed in $\mathbf{H}$}, so $H={\mathbf{H}}$. ∎
@@
(1) By § 3.7, $H$ is definable in $B$. By assumption, $\widehat{H}$ is a definable subgroup of $H$. By Lemma 3.6, the $\widehat{H}-$coset relation on $H$ is smooth; i.e. there exists a Borel map $f$ on $H$ into a Polish space, such that $f(x)=f(y)$ iff $x\widehat{H}=y\widehat{H}$. But by assumption, $\widehat{G}H=G$; so for any $x\in G$ there exists $\hat{x}\in H$ with $\hat{x}\widehat{G}=x\widehat{G}$. Then $f(x^)\hat$$f(\hat{x})$ is well-defined, analytic and hence, Borel, and maps $G$ to a Polish space, with kernel $\widehat{G}$. In other words the $\widehat{G}-$coset equivalence relation is smooth. By Corollary 3.9 $\widehat{G}$ is closed.
@@
So $N$ is an almost strongly minimal structure (contained in the algebraic closure of a strongly minimal definable set), and $N$ has Morley rank $1$. Pick $c\in I$, and let $d$ enumerate $f^{-1}(c)$. Let $A=\mathop{\rm dcl}\nolimits(d)\cap\mathop{\rm acl}\nolimits(0)$, a subset of $N^{eq}$. Then $tp(d/A)$ is stationary. Clearly $A$ does not depend on the specific enumeration $d$. It does not depend on $c$ either, since $tp(c/0)$ is stationary, so $tp(c/0)$ implies a complete type over $\mathop{\rm acl}\nolimits(0)\supseteq A$. Thus $A$ is an $Aut(N)$-invariant subset of $acl(0)$. Since also $A\subset\mathop{\rm dcl}\nolimits(d)$, by either $\aleph_{0}$-categoricity (using finitely many types $tp(d,d^{\prime})$) or $\omega$-stability (using the Morley rank and degree of $tp(d)$), there can be no strictly increasing chain of definable subsets of $A$, so $A$ is essentially finite, i.e. $A=\mathop{\rm dcl}\nolimits(A_{0})$ for some finite $A_{0}\subset A$.
@@
The $H_{c}$-orbit of $d$ is definable in $N$ over $A,c$; so there exists an $A$-definable function $e^{*}$ such that $e^{*}(c)$ codes the $H_{c}$-orbit of $d$. Now $e^{*}$ is $Aut(N/A)$ - invariant, so the finite-index subgroup $Aut(N/A)$ of $\widehat{G}=Aut(N)$ fixes the function $e^{*}$, $Aut(N/A)\leq Aut(N_{0}/e^{*})$; on the other hand $Aut(N_{0}/e^{*},I)=\prod_{i\in I}H_{i}=Aut(N/A,I)$; since the two groups map onto $Sym(I)$ and have the same kernel, the inclusion $Aut(N/A)\leq Aut(N_{0}/e^{*})$ must be an equality $Aut(N/A)=Aut(N_{0}/e^{*})$.
@@
Let $SET_{\leq n}$ be the theory, in the language with one constant symbol, of a nonempty set $\Omega$ with at most $n$ elements; augmented with imaginary sorts coding subsets of $\Omega^{k}$ for $k\leq n$ (i.e. orbits on $\Omega^{n}$ of various subgroups of $Sym(n)$). Thus the models of $SET_{\leq n}$ are finite altogether. (The constant symbol serves to meet the condition $\clubsuit$.)
It is easy to see that $SET_{\leq n}$ eliminates imaginaries to this finite language. Let $SET_{n}$ be the completion where the home sort has exactly $n$ elements.
@@
By Lemma 3.4 (3) $T_{at}$ has a countably saturated model $M=\Omega^{I}$ with $|\Omega|=n$, $|I|=\aleph_{0}$. We have $B=B(M)={\mathop{\operatorname{\rm P}}}(I)$. We observed in Lemma 3.4 (2) that $M$ is internal to $B$, i.e. $M\subset dcl(B,c_{1},\ldots,c_{n})$ for appropriate elements $c_{1},\ldots,c_{n}$. Recalling also the stable embeddedness of $B$ (Proposition 2.3), we have a definable liaison group $H=Aut(M/B)$. (See [13], Appendix B.) Let $G=Aut(M)\cong Sym(I){\ltimes}H$. The same $G$ is also the automorphism group of the countable structure $\coprod_{i\in I}\Omega_{i}$, where $|\Omega_{i}|=|\Omega|$. We use the latter to endow $G$ with a topology of pointwise convergence, and in particular a Borel structure.
@@
Let $e\in M^{eq}$ be an imaginary. Let $A=dcl(e)\cap B^{eq}$. We have $e\in dcl(d_{1},\ldots,d_{m})$ for some $d_{1},\ldots,d_{m}\in M$. By Theorem 2.3, the induced structure on $B$ from $M(d_{1},\ldots,d_{m})$ is the Boolean algebra structure along with constants $\psi(d_{1},\ldots,d_{m})]$ where $\psi$ is an $m$-ary formula of $T$; there are only finitely many such formulas, so the induced structure on $B$ from $(M,e)$ is a reduct of $B$ enriched with finitely many constants; in other words it is the Boolean algebra enriched with an imaginary constant $A$. By weak elimination of imaginaries in $B$ (Proposition 3.5), there exist elements $b_{1},\ldots,b_{k}\in B$ (that we may take to form a Boolean partition of $1$) such that each $b_{i}\in acl(A)$, and $A\subset dcl(b_{1},\ldots,b_{k})$. By stable embeddedness of $B$, $tp(e/A)$ implies $tp(e/B)$.
@@
Now $B$ splits as the product of $k$ Boolean algebras $B_{i}$ (with $1_{B_{i}}=b_{i}$.) Likewise $M$ can be viewed as $M=\prod_{i=1}^{k}M_{i}$ where $M_{i}$ is the part corresponding to $B_{i}$. Using wEI for finite products (Lemma 3.2), $\mathop{\rm acl}\nolimits(e)$ includes imaginaries $e_{i}$ of $M_{i}$, with $e\in\mathop{\rm dcl}\nolimits(e_{1},\cdots,e_{k})$. So we can reduce to the case $k=1,b_{1}=1$. In other words $\mathop{\rm acl}\nolimits(e)\cap dcl(B)=dcl(0)_{B}$, or again, by stable embeddedness, $\widehat{G}:=Aut(M/e)$ maps onto $Aut(B)$. Note also that $\widehat{G}$ is a Borel subgroup of $G$; if $e=a/E$ for a definable equivalence relation $E$ on $M$, we have $g\in\widehat{G}$ iff $g(a)Ea$.
@@
Let $\widehat{H}=Aut(M/B,e)$. Since $M$ is $B$-internal, again by the theory of internality ([13], Appendix B), the liaison group $\widehat{H}$ is an $M$-definable (with parameters) subgroup of $G$, or an intersection of a descending chain of $M$-definable subgroups $\widehat{H}_{j}$. Since $\widehat{H}$ is also definable over finitely many parameters $e$, by $\aleph_{0}$-categoricity of $M$ it is definable; $\widehat{H}=\widehat{H}_{j}$ for some $j$. Viewing $M$ and $G$ as interpretable with parameters over $B$, we see that $\widehat{H}$ is definable with parameters in $B$. By Lemma 3.11, $Aut(M/e)=\widehat{G}$ has a finite index closed subgroup of the form $Aut(M/e^{*})$ for some function $e^{*}$, such that $e^{*}(i)\in\Omega_{i}^{eq}$. Seen as an element of $M$, we have $e^{*}\in\mathop{\rm acl}\nolimits(e)$ (by homogeneity of $M$ and the finiteness of the orbit of $e^{*}$ under $\widehat{G}$). Conversely $Aut(M/e^{*})$ fixes $e$, so $e\in\mathop{\rm dcl}\nolimits(e^{*})$. This proves weak elimination of imaginaries for $(SET_{n})_{at}$. ∎
@@
Now towards the proofs of Theorem 3.1 and the ’single $J$-class’ part of 5.1.
@@ 3.15. Definition of $E^{Z}$
Let $(M,B)\models T_{at}$. Let $D$ be a sort of $T$, and let $Z=Z_{c}\subset D$ be an $M$-definable set in the language of $T_{at}$, with imaginary canonical parameter $c$. Let $C=\mathop{\rm acl}\nolimits(c)\cap M$, i.e. $C$ is the set of elements in ‘real’ sorts of $M$ that are algebraic over $c$.
Our plan is to define equivalence relations $E_{a}^{Z}$ of $T$ on $D$, uniformly for each atom $a$ of $B$. We will show that for some $n$, we have $|D/E_{a}^{Z}|\leq n$ for each atom $a$ of $B$. We will then let $E^{Z}$ be the equivalence relation on $D(M)$, with $xE^{Z}y$ if and only if for all atoms $a$ of $B$, $xE_{a}y$; and show that $Z$ is a union of $E^{Z}$-classes. This will essentially reduce the problem to the case $D/E^{Z}$ is ‘definably pro-finite’, i.e. the Boolean value of $|D/E^{Z}|\leq n$ is $1$, for some $n$.
Let $a\in B,a^{\prime}=1-a$. We can identify $B$ with a product algebra $B_{a}\times B_{a^{\prime}}$, with $a$ identified with $(1,0)$ and $a^{\prime}$ with $(0,1)$. (So $B_{a}=\{y\in B:y\leq a\}$, with $1_{B_{a}}=a$ and $\neg y=a\smallsetminus y$ for $y\in B_{a}$.). Then $M$ also decomposes definably as $M=M_{a}\times M_{a^{\prime}}$, where $M_{a}=M/\Theta_{a}$, $x\Theta_{a}y$ iff $a\leq[x=y]$. In case $M=\prod_{i\in I}M_{i}$, the factor $M_{a}$ can be identified with $\prod_{i\in a}M_{i}$. Note $M_{a},M_{a^{\prime}}$ are interpretable in $(M,a)$, and stably embedded and orthogonal to each other. (The automorphism group of $(M,a)$ is just the product $Aut(M_{a})\times Aut(M_{a^{\prime}})$.) Thus also $D(M)=D(M_{a})\times D(M_{a^{\prime}})$. For $d\in D(M_{a})$, let $Z(d)=\{d^{\prime}\in D(M_{a^{\prime}}):(d,d^{\prime})\in Z\}$.
We obtain from $Z$ alone a definable map $D(M_{a})\to W(M_{a^{\prime}})$ for a certain imaginary sort $W$ mapping $d\in D(M_{a})$ to the code of $Z(d)$. Since $M_{a},M_{a^{\prime}}$ are orthogonal, this map has finite image.  [^1] In general, a definable map from one interpretable set to an orthogonal one must have finite image; this follows from the fact that the graph of the map is a finite union of rectangles. Thus we canonically obtain from $Z$ finitely many pairwise disjoint subsets $Z_{a,1},\ldots,Z_{a,r}$ (for $M_{a}$) and imaginary elements $e_{1},\ldots,e_{r}$ (for $M_{a^{\prime}}$) such that $Z_{i}\mapsto e_{i}$; but the orderings of the $Z_{i}$ and the $e_{i}$ are not canonical.
@@
By elimination of imaginaries in $T$, $E_{a}^{Z}$ is coded by some element $\gamma_{a}$, definable uniformly in an atom $a$ of $B$. By Axioms 2.2(4) of $T_{boole}$, there exists an element $\gamma$ with $\gamma=\gamma_{a}$ (modulo $a$) for each $a$. (Alternatively, this is clearly true in $\prod_{i\in I}M_{i}$ with finite $I$, hence by Lemma 2.9(2) it holds in $M$.) We defined $\gamma$ canonically from $Z$, hence by the definition of $C$ we have $\gamma\in C$.
Define $xE^{Z}y$ if and only if for all atoms $a$ of $B$, $xE_{a}^{Z}y$.
@@
Let $M$ be a sufficiently saturated and homogeneous model of $T_{at}$. Let $Z$ be an $M$-definable subset of $M^{n}$. We need to find a canonical code for $Z$. By EI in $T$, we may view $D/E_{a}$ as a finite subset $F_{a}$ of some sort of $T$. We have $|F_{a}|\leq n$ for some $n$. Using $\clubsuit$, we can throw in a 0-definable element into $F_{a}$. Then $(D/E,B)$ becomes a model of $(SET_{\leq n})_{at}$, definable over $C$, and $\clubsuit$ holds. Let $Z/E^{Z}$ be the image of $Z$ in $D/E^{Z}$; by Lemma 3.17, a code for $Z/E^{Z}$ also serves as code for $Z$. By Lemma 3.4 (2), $D/E^{Z}$ is internal to $B$; since $B$ is a stably embedded sort of $T_{at}$, $D/E^{Z}$ is also stably embedded, and hence $Z/E^{Z}$ can be defined with parameters in $D/E^{Z}$. By Corollary 3.14, $Z/E^{Z}$ can be defined with a canonical parameter; hence so can $Z$.
4. Imaginaries of atomless Boolean powers

### v3

@@ 3.7. Definable groups and automorphism groups
To bridge the gap between Borel and closed, we will use a result of Becker and Kechris. Taking into account the Harrington-Kechris-Louveau theorem [11] or [1] 3.4.3, Theorem 3.4.5 of [1] asserts: Let $\mathbf{X}$} be a Polish space, $\Gamma$ a group acting on $\mathbf{X}$} by homeomorphisms, and let $R$ be the equivalence relation on $\mathbf{X}$} whose classes are the $\Gamma$-orbits. If there is a dense orbit and $R\subset{\mathbf{X}}^{2}$ is meager, then $R$ is not smooth.
@@
If $H$ is meager in $\mathbf{H}$}, it is easy to see that $R$ is meager in ${\mathbf{H}}^{2}$; but this contradicts [1, Theorem 3.4.5]. Hence $H$ is not meager. By [19, Theorem 9.9], $H=HH^{-1}$ contains an open neighborhood $U$ of the identity of $\mathbf{H}$} ; hence $H=HU=\bigcup_{h\in H}hU$ is open in $\mathbf{H}$}, and thus also closed in $\mathbf{H}$}, so $H={\mathbf{H}}$. ∎
@@
(1) By § 3.7, $H$ is definable in $B$. By assumption, $\widehat{H}$ is a definable subgroup of $H$. By Corollary 3.6, the $\widehat{H}-$coset relation on $H$ is smooth, i.e. there exists a Borel map $f$ on $H$ into a Polish space, such that $f(x)=f(y)$ iff $x\widehat{H}=y\widehat{H}$. By assumption, $\widehat{G}H=G$, so for any $x\in G$ there exists $\hat{x}\in H$ with $\hat{x}\widehat{G}=x\widehat{G}$. Then the function $f(x^)\hat$$f(\hat{x})$ is well-defined, analytic and hence, Borel, and maps $G$ to a Polish space, with kernel $\widehat{G}$. In other words the $\widehat{G}-$coset equivalence relation is smooth. By Corollary 3.9 $\widehat{G}$ is closed.
@@
So $N$ is an almost strongly minimal structure (contained in the algebraic closure of a strongly minimal definable set), and $N$ has Morley rank $1$. Pick $c\in I$, and let $d$ enumerate $f^{-1}(c)$. In $N^{eq}$, define the set $A:=\mathop{\rm dcl}\nolimits(d)\cap\mathop{\rm acl}\nolimits(0)$. Then $tp(d/A)$ is stationary. Clearly $A$ does not depend on the specific enumeration $d$. It does not depend on $c$ either, since $tp(c/0)$ is stationary, so $tp(c/0)$ implies a complete type over $\mathop{\rm acl}\nolimits(0)\supseteq A$. Thus $A$ is an $Aut(N)$-invariant subset of $acl(0)$. Since also $A\subset\mathop{\rm dcl}\nolimits(d)$, by either $\aleph_{0}$-categoricity (using finitely many types $tp(d,d^{\prime})$) or $\omega$-stability (using the Morley rank and degree of $tp(d)$), there can be no strictly increasing chain of definable subsets of $A$, so $A$ is essentially finite, i.e. $A=\mathop{\rm dcl}\nolimits(A_{0})$ for some finite $A_{0}\subset A$.
@@
The $H_{c}$-orbit of $d$ is definable in $N$ over $Ac$; so there exists an $A$-definable function $e^{*}$ such that $e^{*}(c)$ codes the $H_{c}$-orbit of $d$. Now $e^{*}$ is $Aut(N/A)$ - invariant, so the finite-index subgroup $Aut(N/A)$ of $\widehat{G}=Aut(N)$ fixes the function $e^{*}$, so $Aut(N/A)\leq Aut(N_{0}/e^{*})$. On the other hand $Aut(N_{0}/e^{*},I)=\prod_{i\in I}H_{i}=Aut(N/A,I)$; since the two groups map onto $Sym(I)$ and have the same kernel, the inclusion $Aut(N/A)\leq Aut(N_{0}/e^{*})$ must be an equality $Aut(N/A)=Aut(N_{0}/e^{*})$.
@@
Let $SET_{\leq n}$ be the theory, in the language with one constant symbol, of a nonempty set $\Omega$ with at most $n$ elements; augmented with imaginary sorts coding subsets of $\Omega^{k}$ for $k\leq n$. Thus the models of $SET_{\leq n}$ are finite altogether. The constant symbol serves to meet the condition $\clubsuit$ (note that for any $k$ there is a 0-definable subset of $\Omega^{k}$.)
It is easy to see that $SET_{\leq n}$ eliminates imaginaries to this finite language (even without the constant symbol). Let $SET_{n}$ be the completion where the home sort has exactly $n$ elements.
@@
By Lemma 3.4 (3) $T_{at}$ has a countably saturated model $M=\Omega^{I}$ with $|\Omega|=n$, $|I|=\aleph_{0}$. We have $B=B(M)={\mathop{\operatorname{\rm P}}}(I)$. We observed in Lemma 3.4 (2) that $M$ is internal to $B$, i.e. $M\subset dcl(B,c_{1},\ldots,c_{n})$ for appropriate elements $c_{1},\ldots,c_{n}$. Recalling also the stable embeddedness of $B$ (Theorem 2.3), we have a definable liaison group $H=Aut(M/B)$ (cf. [14, Appendix B]). Let $G=Aut(M)\cong Sym(I){\ltimes}H$. The same $G$ is also the automorphism group of the countable structure $\coprod_{i\in I}\Omega_{i}$, where $|\Omega_{i}|=|\Omega|$. We use the latter to endow $G$ with a topology of pointwise convergence, and in particular a Borel structure.
@@
Let $e\in M^{eq}$ be an imaginary. Let $A=dcl(e)\cap B^{eq}$. We have $e\in dcl(d_{1},\ldots,d_{m})$ for some $d_{1},\ldots,d_{m}\in M$. By Theorem 2.3, the induced structure on $B$ from $M(d_{1},\ldots,d_{m})$ is the Boolean algebra structure along with constants $\psi(d_{1},\ldots,d_{m})]$ where $\psi$ is an $m$-ary formula of $T$. Since there are only finitely many such formulas, the induced structure on $B$ from $(M,e)$ is a reduct of $B$ enriched with finitely many constants; in other words it is the Boolean algebra enriched with an imaginary constant $A$. By weak elimination of imaginaries in $B$ (Proposition 3.5), there exist elements $b_{1},\ldots,b_{k}\in B$ (that we may take to form a Boolean partition of $1$) such that each $b_{i}\in acl(A)$, and $A\subset dcl(b_{1},\ldots,b_{k})$. By stable embeddedness of $B$, $tp(e/A)$ implies $tp(e/B)$.
@@
Now $B$ splits as the product of $k$ Boolean algebras $B_{i}$ (with $1_{B_{i}}=b_{i}$.) Likewise $M$ can be viewed as $M=\prod_{i=1}^{k}M_{i}$ where $M_{i}$ is the part corresponding to $B_{i}$. Using wEI for finite products (Lemma 3.2), $\mathop{\rm acl}\nolimits(e)$ includes imaginaries $e_{i}$ of $M_{i}$, with $e\in\mathop{\rm dcl}\nolimits(e_{1},\cdots,e_{k})$. So we can reduce to the case $k=1$ and $b_{1}=1$. In other words $A\subset\mathop{\rm dcl}\nolimits(b_{1})=\mathop{\rm dcl}\nolimits(\emptyset)$; so $\mathop{\rm dcl}\nolimits(e)\cap B^{eq}=dcl(0)_{B}$, or again, by stable embeddedness, $\widehat{G}:=Aut(M/e)$ maps onto $Aut(B)$. Note also that $\widehat{G}$ is a Borel subgroup of $G$; if $e=a/E$ for a definable equivalence relation $E$ on $M$, we have $g\in\widehat{G}$ iff $g(a)Ea$.
@@
Let $\widehat{H}=Aut(M/B,e)$. Since by Lemma 3.4(2) $M$ is $B$-internal, again by the theory of internality ([14, Appendix B]), the liaison group $\widehat{H}$ is an $M$-definable (with parameters) subgroup of $G$, or an intersection of a descending chain of $M$-definable subgroups $\widehat{H}_{j}$. Since $\widehat{H}$ is also definable over finitely many parameters $e$, by $\aleph_{0}$-categoricity of $M$ it is definable; $\widehat{H}=\widehat{H}_{j}$ for some $j$. Viewing $M$ and $G$ as interpretable with parameters over $B$, we see that $\widehat{H}$ is definable with parameters in $B$. By Lemma 3.11, $Aut(M/e)=\widehat{G}$ has a finite index closed subgroup of the form $Aut(M/e^{*})$ for some function $e^{*}$, such that $e^{*}(i)\in\Omega_{i}^{eq}$. Seen as an element of $M$, we have $e^{*}\in\mathop{\rm acl}\nolimits(e)$ (by homogeneity of $M$ and the finiteness of the orbit of $e^{*}$ under $\widehat{G}$). Conversely $Aut(M/e^{*})$ fixes $e$, so $e\in\mathop{\rm dcl}\nolimits(e^{*})$. This proves weak elimination of imaginaries for $(SET_{n})_{at}$. ∎
Before proceeding, let us show Lemma 3.13 immediately yields the case of finite sets with arbitrary structure. This will not be used in the sequel since a somewhat more general version will be required, namely enrichments of the infinite product that do not add structure to the Boolean algebra.
@@
We now move towards the proof of weak elimination of imaginaries for the atomic case (cf. Theorem 3.1). The tools that we develop in the following section will play an important role in the mixed case, Theorem 5.1, allowing us to code a subset $Z$ contained in a single $J$-class.
@@ 3.15. Definition of $E^{Z}$
Let $(M,B)\models T_{at}$. Let $D$ be a sort of $T$, and let $Z=Z_{c}\subset D$ be an $M$-definable set in the language of $T_{at}$, with imaginary canonical parameter $c$. Let $C=\mathop{\rm acl}\nolimits(c)\cap M$, i.e. $C$ is the set of elements in ‘real’ sorts of $(M,B)$ that are algebraic over the imaginary parameter $c$. We will write $M_{C}$ for the expansion of $(M,B)$ obtained by naming the elements of $C$.
Our plan is to define equivalence relations $E_{a}^{Z}$ of $T$ on $D$, uniformly for each atom $a$ of $B$. We will show that for some $n$, we have $|D/E_{a}^{Z}|\leq n$ for each atom $a$ of $B$. We will then let $E^{Z}$ be the equivalence relation on $D(M)$, with $xE^{Z}y$ if and only if for all atoms $a$ of $B$, $xE_{a}y$; and show that $Z$ is a union of $E^{Z}$-classes. This will essentially reduce the problem to the case $D$ is ‘definably pro-finite’, i.e. the Boolean value of $|D|\leq n$ is $1$, for some $n$.
Let $a\in B,a^{\prime}=1-a$. We can identify $B$ with a product algebra $B_{a}\times B_{a^{\prime}}$, with $a$ identified with $(1,0)$ and $a^{\prime}$ with $(0,1)$. (So $B_{a}=\{y\in B:y\leq a\}$, with $1_{B_{a}}=a$ and $\neg y=a\smallsetminus y$ for $y\in B_{a}$). Then $M$ also decomposes definably as $M=M_{a}\times M_{a^{\prime}}$, where $M_{a}=M/\Theta_{a}$, $x\Theta_{a}y$ iff $a\leq[x=y]$. In case $M=\prod_{i\in I}M_{i}$, the factor $M_{a}$ can be identified with $\prod_{i\in a}M_{i}$. Note $M_{a},M_{a^{\prime}}$ are interpretable in $(M,a)$, and stably embedded and orthogonal to each other. (Note that the automorphism group of $(M,a)$ is just the product $Aut(M_{a})\times Aut(M_{a^{\prime}})$). Thus also $D(M)=D(M_{a})\times D(M_{a^{\prime}})$. For $d\in D(M_{a})$, let $Z(d)=\{d^{\prime}\in D(M_{a^{\prime}}):(d,d^{\prime})\in Z\}$.
We obtain from $Z$ alone a definable map $D(M_{a})\to W(M_{a^{\prime}})$ for a certain imaginary sort $W$ mapping $d\in D(M_{a})$ to the code of $Z(d)$. Since $M_{a},M_{a^{\prime}}$ are orthogonal, this map has finite image.  [^1] In general, a definable map from one interpretable set to an orthogonal one must have finite image; this follows from the fact that the graph of the map is a finite union of rectangles. Thus we canonically obtain from $Z$ finitely many pairwise disjoint subsets $Z_{a,1},\ldots,Z_{a,r}$ (for $M_{a}$) and imaginary elements $e_{1},\ldots,e_{r}$ (for $M_{a^{\prime}}$) such that $Z_{a,i}\mapsto e_{i}$; but the orderings of the $Z_{a,i}$ and the $e_{i}$ are not canonical.
@@
By elimination of imaginaries in $T$, $E_{a}^{Z}$ is coded by some tuple $\gamma_{a}$ from $M_{a}$, definable uniformly in an atom $a$ of $B$. By Axioms 2.2(4) of $T_{boole}$, there exists an element $\gamma$ with $\gamma=\gamma_{a}$ (modulo $a$) for each $a$. Alternatively, this is clearly true in $\prod_{i\in I}M_{i}$ with finite $I$, hence by Lemma 2.10(1) it holds in $M$. We defined $\gamma$ canonically from $Z$, hence by the definition of $c$ and $C$ we have $\gamma\in C$. So $E_{a}^{Z}$ is definable uniformly in $a$ over $C$.
Define $xE^{Z}y$ if and only if for all atoms $a$ of $B$, $xE_{a}^{Z}y$. So $E^{Z}$ is $C$-definable.
@@
Let $M$ be a sufficiently saturated and homogeneous model of $T_{at}$. Let $Z$ be an $M$-definable subset of $M^{n}$. We need to find a canonical code for $Z$. Recall the definition of $C=\mathop{\rm acl}\nolimits(\ulcorner Z\urcorner)\cap M$ and $M_{C}$ from § 3.15, as well as the equivalence relations $E_{a}$ and $E$.
By EI in $T$, we may view $D/E_{a}$ as a finite subset $F_{a}$ of some sort $S$ of $T$. We have $|F_{a}|<n$ for some $n$. Using $\clubsuit$, we can throw in a 0-definable (in $T$) element $\omega\in S$ into $F_{a}$. Then $D/E$ embeds definably into the sort $S^{M}$ (with its Boolean $T_{at}$-interpreation). Let $Q=(D/E,B)$, with the structure of a model of $(SET_{\leq n})_{at}$: the Boolean algebra structure on $B$, the maps $x=y]:(D/E)^{2}\to B$ and $x=\omega]:(D/E)\to B$. So $Q$ is interpretable in $M_{C}$. Of course, the induced structure on $Q$ from $M_{C}$ may be richer; but we have:
Claim .   Let $W\subset Q^{k}$ be $M$-definable. Then $W$ is definable with parameters in $Q$.
Proof.
By Lemma 3.4 (2), $D/E^{Z}$ is internal to $B$; with parameters from itself. The statement thus reduces to subsets of $B^{l}$, and follows from embeddedness and stable embeddedness of $B$. ∎
In particular, $Z/E^{Z}$ is definable with a parameter $c^{\prime}$ from $Q^{l}$. By Corollary 3.14, $Z/E^{Z}$ can be defined with a weakly canonical parameter $c^{\prime}$ from $Q^{l}$; i.e. $c^{\prime}\in\mathop{\rm acl}\nolimits(\ulcorner Z/E^{Z}\urcorner)$ Now $c^{\prime}$ is equi-definable with an element $c^{\prime\prime}$ of $S^{M}$. So $c^{\prime\prime}\in C$. By Lemma 3.17, $Z$ is $\ulcorner Z/E^{Z}\urcorner$-definable, hence $C$-definable.
4. Imaginaries of atomless Boolean-valued models